# TRIM: Token Reduction and Inference Modeling for Cost-Effective Language Generation

### v3

@@ Abstract
The inference cost of Large Language Models (LLMs) is a significant challenge due to their computational demands, specially on tasks requiring long outputs. However, natural language often contains redundancy, which presents an opportunity for optimization. We have observed that LLMs can generate distilled language—concise outputs that retain essential meaning, when prompted appropriately. We propose TRIM, a pipeline for saving computational cost in which a shorter distilled output from the LLM is reconstructed into a full narrative by a smaller model with lower inference costs. Our experiments show promising results, particularly in general knowledge domains with 20.58% saved tokens on average with tiny decrease in evaluation metrics, hinting that this approach can effectively balance efficiency and accuracy in language processing tasks.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities across a wide range of tasks, from natural language understanding to creative content generation. However, the computational cost of inference and the associated energy consumption present significant challenges. As the demand for AI applications continues to grow, these costs are expected to escalate, raising concerns about sustainability and accessibility Wu et al. (2022). Addressing these issues is crucial to ensure that future advancements in AI remain feasible and widely deployable. Hence, we propose a new technique to save tokens at inference time skipping less informative and semantic-irrelevant function words. Later on, we propose to regenerate them based on the context with a cost-effective alternative, leveraging on the feature that natural language is considerable redundant.
As part of our feasibility analysis, we empirically verified that this task is achievable by state-of-the-art LLMs, even though they are not specifically designed for such tasks, and the altered output may not have been encountered in any part of the training corpus. Given that LLMs can generate this distilled output, the omitted words can be reinserted into the output text using a language model with a low inference cost. Studies in cognitive linguistics suggest that human readers naturally infer missing function words based on context Ferreira and Patson (2007). Similarly, transformer-based models use masked language modeling, where parts of a sentence are intentionally masked, and the model learns to predict the masked tokens based on context Devlin et al. (2019). Our hypothesis is that small language models can mirror this human reading abilities of filling the gaps for low-information inferable terms.
In this work, the main contribution is the proposal of a novel pipeline for reducing the inference costs of LLMs through the suppression and later-on reconstruction of function words. This approach is orthogonal to other optimization techniques, and could be applicable as LLMs continue to grow in size and capabilities. We also propose an algorithm to check and define the applicability of this technique in different domains, selecting the most proper function words set, and analyzing the lose in performance as the percentage of saved tokens increases. Additionally, we provide an experimental evaluation in the context of general knowledge question-answering. In this setting, the output length is more relevant than the input query, making the applicability of our technique more evident. The results show that even suppressing around one fifth of the full text, models that are hundreds of times smaller can reconstruct distilled text without losing significant text quality.
@@ 2.1 Prompt Compression
One alternative to decrease the computational cost is setting the focus on reducing the input by prompt compression. This could be a token-level removal of less relevant words Jiang et al. (2023); Pan et al. (2024); Fu et al. (2024), or more elaborated approaches such as context-aware sentence encoding Liskavets et al. (2024). This task operates on the input rather than on the output. Therefore, our approach would be complementary to prompt compression. Moverover, prompt compression does not require a later reconstruction of original input text, as it pertains to the encoding step and the result of the compression is handled internally with vectorized representations.
@@ 2.2 Inference Optimizations
For the generative-LLM inference, two phases are recognizable. In the first one, all the input tokens from the prompt are computed in parallel. In the second one, tokens are generated in successive passes using previously generated tokens, and the key-value cache collected in the first phase Miao et al. (2023a). Given the memory-intense requirement, the token generative phase might be allocated en specialized hardware Patel et al. (2024). On the other hand, smaller LLM models can be used to generate multiple output sequences in parallel that then are merged and verified Miao et al. (2023b). Another way to save costs at inference time consists of building a sequence of increasing-size LLMs, where a (larger) model at step $N$ is only queried when the answer of the model at step $N-1$ is not deem reliable Chen et al. (2023).
@@ 2.2.1 Text Reconstruction
One of the core learning ideas in Natural Language Processing (NLP) is having a vectorized representation that allows models to reconstruct back the original text Oshri and Khandwala (2015). Nevertheless, NLP tasks, such as masking Devlin et al. (2019) or in a more general scope, infilling Donahue et al. (2020), rely on markers indicating where to substitute the text. In our approach, the input text at inference time would not have indication of what positions contained an omitted word. Text with omitted words can also be used as negative examples, to improve the performance of automatic translators Yang et al. (2019), but to our knowledge omitted/original texts have not been used to train for inference efficiencies.
@@ 3 Exploratory Research
The feasibility of our approach relies on the LLM being able to perform the task of generating an answer omitting a set of terms that do not contribute to the semantic meaning. This is a task for which the LLM was not originally trained. To verify this assumption, we performed an exploratory analysis where we used the Gpt-4o model OpenAI (2024), and developed a prompt to carry out this task indicating also that omitted terms should be easily inferrable in a later stage. The prompt includes a pre-defined set of terms composed by determiners and conjunctions with a total of 23 terms such as: (the, a, an, and, but, so, that). See supplementary material (Section A.1) for details. We asked the LLM to generate the answer to 500 knowledge questions (See dataset in Section 6) with and without the previously defined prompt. We computed the average number of terms from the set present in both types of answers (Table 1). We observed how the LLM is able to complete the task effectively, generating the answer omitting nearly all terms. Additionally, the average is not zero, indicating that in some scenarios the LLM preserves some of these terms pro-actively deciding that they are not easily reconstructible a posteriori in that scenario.
@@
Table 1: Average count of the terms present in the text.
4 Semantic Relevance Distribution Of Words
The importance of some terms when measuring the semantic meaning of texts is a well-documented phenomenon, particularly in English. A study by Zipf (1949) highlighted that "speakers tend to use more common words more frequently to reduce the cognitive and physical effort involved in communicatingZipf (1949). Certain words, such as articles and determiners, often contribute little to the overall semantic meaning of a sentence. These words are frequently classified as function words, which serve a grammatical purpose rather than adding significant content or meaning to a sentence Jurafsky and Martin (2000); Manning and Schütze (1999).
Even though function words provide little semantic meaning, not all of them can be effectively reconstructed on the basis of context. Some of them, such as pronouns or quantifiers, cannot be always inferred based on context. For example, in the sentence "I went to the marathon in the city center", if we remove the function words, the sentence will be "went marathon city center". The terms "the", "to", and "in" are easily recovered based on the context, but the first pronoun could be any of the personal pronouns.
The categorization and role of function words are clear from the English grammar and linguistic theory. However, few NLP studies have analyzed them regarding how they affect representation learning Kim et al. (2019). Besides, our interest in this paper lies on how easy it is for a language model to infer the function words, so then we can pre-select the best set of words to include in the instruction prompt. For this purpose, we propose a new algorithm based on the use of language models trained with the masking technique that allows to obtain a sorted list of the most easily inferable words based on the context. The algorithm focuses on calculating the difference in probability between the actual word and the next most probable word in a fixed position based on the context. This difference represents the prevalence of the actual word over the distribution of the rest of possible words. This means that, the bigger the difference, the less doubt a language model has about choosing that term over the rest, and, therefore, it is easily inferable based on the context.
For a given text $c$ (e.g. a paragraph from a corpus) and a word $w$, let $c^{\prime}$ be the text $c$ with the term $t$ masked. We define the probability difference as:
@@ 4 Semantic Relevance Distribution Of Words
where $P(w_{\rm{alt}}\mid c)$ is the probability of the most probable alternative term ($w_{\rm{alt}}\neq w$) given the same context, obtained using the LM. The sorted list of function words based on this criteria is computed following Algorithm 1. After computing $\Delta P$ for all function words in each corpus fragment, it computes the average for each word, and returns the list, sorted by this average.
Algorithm 1 Terms sorted by ease of inference
1:Input: Word set $T=\{w_{1},w_{2},\dots,w_{n}\}$,
@@
3:   language model $LM$
4:Output: Sorted T by probability difference $\Delta P_{j}$
@@
6:     for each $t_{j}\in T$ do
7:         if $t_{j}\in c_{i}$ then
8:              Compute $\Delta P_{i,j}$ with $LM$
@@
15:$T^{\prime}$ = Sort $T$ by $\Delta P_{j}$ in descending order
16:return T’
@@ 5 Cost-Saving Pipeline
The exploratory research showed the feasibility of the initial idea. Now, we cover the development and specification of the approach. We propose TRIM (Token Reduction and Inference Modeling pipeline for cost-effective language generation), a pipeline for wrapping the use of generative LLMs (both closed and open) to save costs during output generation through the philosophy of omitting the generation of inferable function words, and reconstructing the full text using a much smaller and optimized language model later on. This pipeline can be seen in Figure 1. It can be defined by the following elements:
Inferable terms: Let $S$ be a set of words deemed as easily inferable based of the context (i.e. stop words, functions words…). This set can be experimentally collected from a corpus using Algorithm 1 proposed earlier.
Generation model: Let $M_{G}$ be the generative LLM model capable of generating a response to a query, omitting terms from $S$. The response is called distilled answer $A_{D}$.
Reconstructor model: Let $M_{R}$ be an encoder-decoder model fine-tuned to reconstruct truncated output into a full text with regards of the specific set of terms $S$.
Instruction Prompt: Let $\mathcal{P}$ be the prompt designed to instruct the LLM in the task of omitting words given $S$.
Being $Q$ the query from the user and $A$ the expected answer returned to the user in normal conditions, the pipeline process could be exemplified as:
$A_{D}=M_{G}(Q,P(S))$
@@ 6 Datasets & Models
For the experimentation, we used two datasets. First, the Wikipedia dump, which is composed of a large amount of well-written and grammatically correct data on the domain of knowledge Foundation . Different subsets from it will be used for analyzing the ease of inferring functions words, and also for training and testing reconstruction models.
Second, we used a synthetic knowledge question-answering dataset we have created for this project. We required an evaluation that aligns with the proposed TRIM pipeline where the reconstruction model reconstructs distilled text generated by an LLM. For this purpose, we generated 9,000 knowledge questions on different topics. For the generation, we took 30 general knowledge topics, 20 concepts on each of them and 5 keywords associated to each concept, creating various questions from templates regarding explanation, summarization and concept comparison. See supplementary material (Section A.2) for specifications, and an example of these questions.
Later, for each question, GPT-4o was requested to generate both the answer and the distilled version of the answer. To achieve this, we developed the instruction prompt $\mathcal{P}$ that compacts this request with the desired term set $S$, following the few-shot philosophy. Ideally, the generation of the original and distilled answer should be requested in different sessions, so the distilled one would simulate the real scenario of our pipeline, and the original would be the ground truth. Otherwise, the LLM would have the context of the original to generate the distilled version. However, after carrying some experiments we realized that doing it in different sessions produce additional variability (e.g., synonyms, grammatical alternative and paraphrasing) that prevent us to use them in the evaluation (see Supplementary Material A.3). Under the assumption that the LLM can generate distilled text directly as shown in Section 3, we keep the single session generation to focus the evaluation on the ability of reconstructing the distilled narrative. Thus, this second dataset called from now on DKA dataset - Distilled Knowledge Answers, is composed of distilled answers with their original answers that only differ on omissions from the input $S$.
Regarding algorithm 1, in charge of obtaining the sorted set of easily inferable words based on the context, the RoBERTa model developed by Facebook AI has been used Liu et al. (2019), a lightweight and robust optimized BERT model that allows us to obtain the probabilities of occurrence for each of the terms in a specific masked position.
@@
On the side of the reconstructor model $M_{R}$, we consider (1) BART, Facebook AI’s model recognized as one of the early transformer-based models with an auto-regressive decoder for text generation Lewis et al. (2019); (2) T5 from Google, which has demonstrated strong performance in summarization and translation tasks Raffel et al. (2023) and finally, (3) Qwen2.5 model from Alibaba, specifically the 0.5B and 1.5B variants, as it was positioned among the top instruct model at its size at the time of comparison, according to the Open LLM Leaderboard Yang et al. (2024) Fourrier et al. (2024). All these models are at least two orders of magnitude smaller than state-of-the-art LLMs, which makes them good candidates to achieve the reconstruction with lower computational costs. Table 2 depicts model size in billion parameters for these models.
@@ 7 Evaluation
The evaluation is divided into two experiments. The first one applies Algorithm 1 for the identification of a good set of inferable terms $S$ without notable downgrade in the metrics. In the second experiment, once the set $S$ has been fixed, several language models will be trained and evaluated in the task of the reconstructor model $M_{R}$. We also analyzed if the results validate the applicability of the complete pipeline without noticing relevant deterioration in the output.
For the evaluation, we thought it was necessary to find a specific metric that would evaluate the recovery of the terms from the inferable set. The common NLP metrics focus on comparing two pieces of text that can come in different sizes (i.e., word, sentence, paragraph…), but they do not offer the versatility of focusing only on a specific set of terms according to their exact position, covering the cases of omission, substitution, matching or insertion. The Needleman-Wunsch algorithm, used in bio-informatics to align protein or nucleotide sequences, offers these desired features Needleman and Wunsch (1970), so we adapted it to align the paragraphs of the original and reconstructed text to effectively maintain the alignment over scenarios where there are missing or extra terms. After applying the algorithm, we can compare each of the positions in the alignment and cover the cases mentioned above, obtaining precision, recall and F1 values with regard to the inferable set. From now on, these metrics will be referred as $\Theta$ metrics.
@@
The rest of evaluation metrics employed compare the whole texts and encompass three primary aspects: grammatical similarity, semantic similarity, and coherence. Grammatical similarity between two texts is assessed using the metrics SacreBLEU, METEOR, and ROUGE, which primarily perform n-gram-based comparisons Post (2018); Banerjee and Lavie (2005); Lin (2004). For semantic similarity, we use cosine similarity, which measures the similarity between two texts using RoBERTa embeddings Liu et al. (2019); Salton et al. (1975). To evaluate textual coherence, we utilize the perplexity metric, which quantifies coherence and fluency based on how easily a language model (in our case, GPT-2 Radford et al. (2019)) can predict the next tokens, where lower values indicates higher fluency and coherence Manning and Schütze (1999). These experiments were performed on an AWS "g4dn.xlarge" machine with an Nvidia "T4" GPU with 16GB of graphics memory.
7.1 Inferable terms set

### v5

@@ Abstract
The high inference cost of Large Language Models (LLMs) poses challenges, especially for tasks requiring lengthy outputs. However, natural language often contains redundancy, which presents an opportunity for optimization. We have observed that LLMs can generate distilled language (i.e., concise outputs that retain essential meaning) when prompted appropriately. We propose TRIM, a pipeline for saving computational cost in which the LLM omits a predefined set of semantically irrelevant and easily inferable words based on the context during inference. Then, a specifically trained smaller language model with lower inference cost reconstructs the distilled answer into the ideal answer. Our experiments show promising results, particularly on the proposed NaLDA evaluation dataset focused on the reconstruction task, with 19.4% saved tokens on average for GPT-4o and only a tiny decrease in evaluation metrics. This suggests that the approach can effectively balance efficiency and accuracy in language processing tasks.
@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities across a wide range of tasks, from natural language understanding to creative content generation. However, the computational cost of inference and the associated energy consumption present significant challenges. As the demand for AI applications continues to grow, these costs are expected to escalate, raising concerns about sustainability and accessibility Wu et al. (2022). Addressing these issues is crucial to ensure that future advancements in AI remain feasible and widely deployable. Hence, we propose a new technique to save tokens at inference time by skipping less informative function words that are easily-inferable based on the context. Later on, we propose to regenerate them based on the context with a cost-effective alternative, leveraging the fact that natural language is considerably redundant.
As part of our feasibility analysis, we verified that state-of-the-art LLMs can omit a predefined set of easily-inferable words, even without being specifically designed for this task. Studies in cognitive linguistics suggest that human readers naturally infer missing function words based on context Ferreira and Patson (2007). Similarly, transformer-based models use masked language modeling to predict the intentionally masked tokens based on the context Devlin et al. (2019). Our hypothesis is that small language models can mirror this human reading abilities of filling the gaps for low-information easily-inferable words.
In this work, the main contribution is the proposal of a novel pipeline called TRIM. This pipeline reduces the inference cost of LLMs by omitting easily-inferable words, which are later on reconstructed using much smaller specifically trained language models. This approach is orthogonal to other optimization techniques, and could be applicable as LLMs continue to grow in size and capabilities. To support this pipeline, we propose also an algorithm designed to sort a list of words by how easily they can be inferred from the context based on a corpus of texts, as well as a technique to experimentally define the length of the set of words based on metrics losses as the percentage of saved tokens increases. To support our evaluation, we developed a new dataset specifically designed for the reconstruction task and covering five different text generation types. Additionally, we introduce a new evaluation metric tailored to assess reconstruction performance, focusing on the words omitted and subsequently inferred. The results show that even omitting around one fifth of the full text, models that are hundreds of times smaller can reconstruct distilled text without losing significant text quality according to different metrics, keeping the semantic meaning practically intact.
@@ 2.1 Prompt Compression
One alternative to decrease the computational cost is setting the focus on reducing the input by prompt compression. This could be a token-level removal of less relevant words Jiang et al. (2023); Pan et al. (2024); Fu et al. (2024), a learned policy guided by reinforcement learning Jung and Kim (2024), context-aware sentence encoding approached Liskavets et al. (2024), and more Zhang et al. (2025). These techniques operate on the input rather than the output, making our approach complementary to prompt compression.
@@ 2.2 Inference Optimizations
The research of general inference optimization techniques around LLM has been a hot topic with several different techniques proposed. Some of them relate to the computation of the tokens from the prompt using parallel and key-value cached techniques Miao et al. (2023a); Patel et al. (2024). In other implementations, smaller LLM models can be used to generate multiple output sequences in parallel which are then merged and verified Miao et al. (2023b). Another way to save costs at inference time for complex tasks consists of building a sequence of increasing-size LLMs, where a (larger) model at step $N$ is only queried when the answer of the model at step N1N-1 is not deem reliable Chen et al. (2023). Furthermore, several techniques around the instance and cluster level optimization have been proposed Zhen et al. (2025). All of these techniques are compatible with our approach as they tackle the problem from different points of view, making them orthogonal to our proposal.
@@ 2.2.1 Text Reconstruction
Text reconstruction has been studied since long time ago. Text with mutilated characters require reconstruction Miller and Friedman (1957) due to errors that occur in message transmission or typing. One of the core learning ideas in Natural Language Processing (NLP) is having a vectorized representation that allows models to reconstruct back the original text Oshri and Khandwala (2015). Nevertheless, NLP tasks, such as masking Devlin et al. (2019) or in a more general scope, infilling Donahue et al. (2020), rely on markers indicating where to substitute the text. In our approach, the input text at inference time would not have indication of what positions contained an omitted word. Text with omitted words can also be used as negative examples, to improve the performance of automatic translators Yang et al. (2019), but to our knowledge omitted-content texts have not been used to train for inference efficiencies purposes.
@@ 3 Exploratory Research
The feasibility of our approach relies on the LLM being able to perform the task of generating an answer omitting a set of words that do not contribute to the semantic meaning. This is a task for which the LLM was not originally trained. To verify this assumption, we performed an exploratory analysis where we used the GPT-4o model from OpenAI OpenAI (2025), and developed a prompt to carry out this task indicating also that omitted words should be easily inferrable in a later stage. The prompt includes a pre-defined set of words composed by determiners and conjunctions with a total of 23 words such as: (the, a, an, and, but, so, that). See supplementary material (Section A.1) for details regarding the full set of words, designed prompt and further analysis. We took 500 knowledge questions (See dataset in Section 6) and gave them to the LLM in two rounds, first with the previously defined prompt appended and then without it to collect the original expected answer. We computed the average number of words from the set present in both types of answers (Table 1). We observed how the LLM is able to complete the task effectively, generating the answer omitting nearly all words. Interestingly, the average is not zero, indicating that in some scenarios the LLM preserves some of these words pro-actively deciding that they are not easily-inferable a posteriori based on the context in that scenario.
@@
Table 1: Average count of the words present in the text.
4 Semantic Relevance Of Words
The importance of certain words in conveying semantic meaning is a well-documented phenomenon in English Zipf (1949). Common words, such as articles and determiners, often contribute little to the overall meaning of a sentence. These are typically classified as function words, serving a grammatical role rather than adding substantial content or meaning. Jurafsky and Martin (2000); Manning and Schütze (1999). Even though function words provide little semantic meaning, not all of them can be effectively reconstructed on the basis of context. Some of them, such as pronouns or quantifiers, cannot be always inferred based on context. For example, in the sentence "I went to the marathon in the city center", if we remove the function words, the sentence will be "went marathon city center". The words "the", "to", and "in" are easily recovered based on the context, but the first pronoun could be any of the personal pronouns.
The categorization and role of function words are clear from the English grammar and linguistic theory. However, few NLP studies have analyzed them regarding how they affect representation learning Kim et al. (2019). Our interest in this paper lies on how easy it is for a language model to infer the function words based on the context, so then we can select a set of easily-inferable function words to be included in the instruction prompt. For this purpose, we propose a new algorithm that leverages language models trained with masking technique. This approach enables the sorting of an input word list (i.e. function words, stop words…) based on how easily are them inferred based on the context The algorithm focuses on calculating the difference in probability between the actual word and the alternative most probable word in a fixed position based on the context. This difference represents the prevalence of the actual word over the distribution of the rest of possible words. This means that, the bigger the difference, the less doubt a language model has about choosing that word over the rest, and, therefore, it is easily-inferable based on the context.
For a given text $c$ (e.g. a paragraph from a corpus) and a word $w$, let $c^{\prime}$ be the text $c$ with the word $w$ masked. We define the probability difference as:
@@ 4 Semantic Relevance Of Words
where $P(w_{\rm{alt}}\mid c^{\prime})$ is the probability of occurrence of the most probable alternative word ($w_{\rm{alt}}\neq w$) given the same context, obtained using the LM. The sorted list of function words based on this criteria is computed following Algorithm 1. After computing $\Delta P$ for all function words in each corpus fragment, it computes the average for each word, and returns the input list of words sorted by this average.
Algorithm 1 Words sorted by ease of inference
1:Input: Word set $W=\{w_{1},w_{2},\dots,w_{n}\}$,
@@
3:   language model LMLM
4:Output: Sorted W’ by probability difference $\Delta P_{j}$
@@
6:  for each $w_{j}\in W$ do
7:   if $w_{j}\in c_{i}$ then
8:     Compute $\Delta P_{i,j}$ with LMLM
@@
[…]
16:return W’
@@ 5 Cost-Saving Pipeline
The exploratory research showed the feasibility of the initial idea. Now, we cover the development and specification of the pipeline. We propose TRIM (Token Reduction and Inference Modeling pipeline for cost-effective language generation), a pipeline for wrapping the use of generative LLMs (both closed and open) to save costs during output generation of natural language through the philosophy of omitting the generation of easily-inferable low-informative words. Later on, a much smaller and optimized language model handles the reconstruction of the full text. The pipeline, also depicted in Figure 1, comprises the following elements:
Instruction: Let $I$ be the instruction where the task of omitting easily-inferable words based on a pre-defined set is explained in detail to the LLM.
Easily-Inferable Word Set: Let $W$ be a set of words with low semantic relevance (i.e. function words, stop words…) deemed as easily-inferable based of the context. This set can be experimentally collected from a corpus using Algorithm 1.
Instruction Prompt: Let $\mathcal{P}$ be the prompt designed to instruct the LLM in the task of omitting words given $W$ based o the instruction $I$.
Generation Model: Let $M_{G}$ be the generative LLM model capable of generating a response to a query, omitting words from $W$. The response is called distilled answer $A_{D}$.
Reconstruction Model: Let $M_{R}$ be an encoder-decoder model fine-tuned to reconstruct truncated output into full narrative text with regards of the specific set of words $W$.
Being $Q$ the user query and $A$ the expected answer returned to the user in normal conditions, the pipeline process could be exemplified as:
$A_{D}=M_{G}(Q,P(W))$
@@ 6 Datasets & Models
For the experimentation, we used two datasets. Firstly, the Wikipedia dump, which is composed of a large amount of well-written and grammatically correct data on the domain of knowledge Foundation . Different subsets from it will be used for analyzing the ease of inferring functions words, and also for training and testing reconstruction models.
Secondly, we have created a new evaluation dataset specifically tailored to evaluate the performance of the reconstruction model $M_{R}$ in the context of the TRIM pipeline where it reconstructs distilled text generated by a given generation model $M_{G}$, resembling a real use case scenario. In this dataset, called Natural Language Distilled Answers dataset - NaLDA, we cover five main task types of the natural language generation field: create, knowledge-answer, re-write, summarize and translate. The selection of this 5 task types comes from the research in the literature on natural language generation tasks performed by LLMs, like the building of the Alpaca Dataset Taori et al. (2023). The NaLDA dataset is composed of 44.800 entries across the 5 task types, where each entry instructs a given generation model $M_{G}$ to generate, for a given word set $W$, a distilled answer $A_{D}$ and the ideal answer $A$ to a task. With these two answers, a reconstruction model $M_{R}$ can be later on evaluated. See supplementary material (Section A.2) for specifications on the NaLDA dataset elaboration, prompts and examples.
For Algorithm 1, which generates a sorted set of contextually easily-inferable words, we use the RoBERTa model Liu et al. (2019). RoBERTa efficiently computes word occurrence probabilities at masked positions based on context.
As the generation model $M_{G}$ for the evaluation, we consider models from (1) OpenAI OpenAI (2025), such as GPT-4o (gpt-4o-2024-08-06), GPT-4.1-mini (gpt-4.1-mini-2025-04-14) and GPT-4.1 (gpt-4.1-2025-04-14) and models from (2) Anthropic Anthropic (2025), including Claude Sonnet 4 (claude-sonnet-4-20250514) and Claude Opus 4.1 (claude-opus-4-1-20250805).
@@
To evaluate reconstruction models, we trained the following models: (1) BART from Facebook AI, an early transformer with an auto-regressive decoder Lewis et al. (2019); (2) Google’s T5, known for strong summarization and translation Raffel et al. (2023); and (3) Alibaba’s Qwen2.5 (0.5B and 1.5B variants), top-ranked among instruct models of similar size according to the Open LLM Leaderboard Yang et al. (2024); Fourrier et al. (2024). These models are much smaller than state-of-the-art LLMs, making them suitable for reconstruction tasks with lower computational costs. Table 2 shows their parameter sizes.
@@ 7 Evaluation
The evaluation is divided into two experiments. The first one applies Algorithm 1 for the identification of a good set of easily-inferable words $W$ without notable downgrade in the metrics. In the second experiment, once the set $W$ has been fixed, several language models will be trained and evaluated in the task of the reconstruction model $M_{R}$. We also analyzed if the results validate the applicability of the complete pipeline without noticing relevant deterioration in the output.
@@
The general evaluation metrics employed to compare texts encompass three primary aspects: grammatical similarity, semantic similarity, and coherence. Grammatical similarity between two texts is assessed using SacreBLEU, METEOR, and ROUGE, which primarily perform n-gram-based comparisons Post (2018); Banerjee and Lavie (2005); Lin (2004). For semantic similarity, we use cosine similarity, which measures the similarity between two texts using RoBERTa embeddings Liu et al. (2019); Salton et al. (1975). For textual coherence, we use the perplexity metric, which quantifies coherence and fluency based on how easily a language model (in our case, GPT-2 Radford et al. (2019)) can predict the next tokens, where lower values indicates higher fluency and coherence Manning and Schütze (1999). These experiments were performed on AWS "g4dn.xlarge" machines with Nvidia "T4" GPU with 16GB of memory.
7.1 Metrics for Reconstruction Task
Furthermore, we thought it was necessary to find a specific metric that would evaluate the reconstruction of the words from the easily-inferable word set $W$. The general NLP metrics focus on comparing two pieces of text that can come in different sizes (i.e., word, sentence, paragraph…), but they do not offer the versatility of focusing only on a specific set of words according to their exact position in the ideal text, covering the cases of omission, substitution, match or insertion. To this extent, we developed a new evaluation metric specifically focused on evaluating the performance of the reconstruction task, given the ideal text and the reconstructed one. We first had to find a way of matching the two texts, so we can later on check each word position. The Needleman-Wunsch algorithm, used in bio-informatics to align protein or nucleotide sequences, offers this desired feature of alignment Needleman and Wunsch (1970). Then, we analyzed and adapted it to our scenario, effectively aligning two texts (one being the original and one being the altered one) even over scenarios where there are missing, different or extra words.
After applying the algorithm, we can compare each positions in the alignment. Since we know beforehand the set of easily-inferable words $W$ that have been omitted, we can account for scenarios involving omission, substitution, match or insertion for those words. With this information, we obtain precision, recall and F1 values to get a clearer view of the quality of the reconstruction task. We consider as true positives (TP) the match scenario, as false positive (FP) scenarios of inferring a word that is incorrect (insertion and substitution), and as false negative (FN) scenarios where the ideal word is not inferred (omission and substitution). From now on, these metrics will be referred as $\Theta$ metrics.
7.2 Inferable Word Set